OpenStudy (anonymous):

A particle moves along the x axis in such a way that its x axis varies with time t according to the relation x=2-5t+6t^2. Find the initial velocity.

OpenStudy (abhisar):

\(\sf velcoity = \frac{dx}{dt}\\ V(t)=\frac{\huge d}{\huge dt}(2-5t+6t^2)=-5+12t\)

OpenStudy (abhisar):

Initial velocity is the velocity at t=0 Plug in the value of t as 0 i the equation v=-5+12t and find the value of v